# From Semantics, Scene to Instance-awareness: Distilling Foundation Model for Open-vocabulary Situation Recognition

### v1

@@ 1. Introduction
The ability to conduct situation recognition in the scene is one of the essential roles of vision (Dai et al., 2023; Liu et al., 2023a; Zhang et al., 2025) and language (Touvron et al., 2023; Li et al., 2023b) research, with broad applications in assistive technology such as autonomous driving (Pan et al., 2024; Liao et al., 2024; Wang et al., 2024a) and visual impairments (Zhao et al., 2024a). Recently, Multimodal Large Language Models (MLLMs) (Dai et al., 2023; Zhu et al., 2023a) have demonstrated remarkable zero-shot scene understanding and can be applied across various domains (Gao et al., 2025; Cai et al., 2024a; Li et al., 2023a). However, many of them rely on smaller large language model (LLM) counterparts (e.g., InstructBLIP (Dai et al., 2023), TinyLLaVA (Liu et al., 2024)), which often underperform in tasks like Grounded Situation Recognition (GSR) (Pratt et al., 2020; Jiang and Riloff, 2023) that require deep comprehension (Wei et al., 2022b; Lu et al., 2024), as reflected by the Top-1 activity prediction accuracy (Top-1-all-verb) in Figure 2. Moreover, although these models are relatively small compared to much larger ones (e.g., with 34B (Young et al., 2024; Touvron et al., 2023) parameters), fine-tuning and deploying such massive models for GSR remain challenging due to their substantial computational and resource demands. This issue is particularly critical for apply GSR to many assistive technologies, which often depend on low-resource edge devices rather than heavy servers with modern GPUs (Cai et al., 2024b; Wang et al., 2023; Mittal, 2024). Addressing this problem is crucial for advancing the development of small and efficient GSR model capable of accurately interpreting complex scenes while preserving the generalization capabilities of MLLMs, potentially benefiting a wide range of assistive technologies.
@@ 2.2. Knowledge Distillation of Large Models
Recent advanced works distill large model capabilities into smaller ones (Zhang et al., 2024; Zhao et al., 2024c; Sun et al., 2023; Wei et al., 2024), demonstrating promising results. Minmax (Wang et al., 2025) formulated dataset distillation as a minmax optimization problem and proposed neural characteristic function discrepancy to effectively measure distributional differences, enabling compact and high-quality synthetic dataset generation. PRR (Zhao et al., 2024c) introduces a retrieval-based Chain-of-Thought (CoT) (Wei et al., 2022c) distillation technique, which transfer knowledge from LLMs to smaller language models, enhancing the performance of the question answering tasks. Tinyllm (Tian et al., 2024) introduces a new knowledge distillation paradigm, where a small student LLM learns from multiple large teacher models, effectively capturing knowledge from multiple rationals while maintaining efficiency. Some methods focus on knowledge distillation through efficient prompt-tuning (Jia et al., 2022a; Liu et al., 2025; Bahng et al., 2022; Chen et al., 2023). PromptMM (Wei et al., 2024) enhances recommender systems by leveraging prompt-tuning, enabling efficient distillation to bridge the semantic gap across multi-modal contexts. PromptKD (Li et al., 2024) leverages soft prompt-based imitation on unlabeled domain images, allowing a lightweight target model to acquire knowledge from a large teacher model through a novel unsupervised distillation approach. Differently, this work encapsulates the information of rich rationales and visual prompts from the teacher model to effectively distill multimodal knowledge into a student model for improved Ov-GSR.
@@ 2.3. Open-vocabulary Tasks
The most similar work to Ov-GSR is end-to-end open-vocabulary HOI detection (Wang et al., 2022; Lei et al., 2024c), which simultaneously recognizes actions in an image while detecting humans and objects. THID (Wang et al., 2022) distills and utilizes transferable knowledge from the pretrained CLIP model, integrating multimodal features into a joint visual-text space to enhance open-vocabulary interaction detection. CMD-SE (Lei et al., 2024c) presents a novel HOI detection framework that distills fine-grained human body part semantic knowledge from LLM to enhance interaction recognition. The proposed Ov-GSR focuses on recognizing the activity first, followed by the detection of multiple entities within an image. In contrast to CMD-SE, our model adopts a different approach by aligning multimodal knowledge from semantics, scene to instance-level, effectively distilling this knowledge to bridge the gap between seen and unseen situations while enhancing rare situation awareness.
@@ 3.1. Problem Overview and Motivation
Distill Knowledge from Large Models with Prompts: Knowledge distillation (Xu et al., 2024; Gu et al., 2024a; Ning et al., 2023) has emerged as a key technique to alleviate the substantial computational demands of modern MLLMs by training smaller student models to replicate the behavior of larger teacher models, significantly reducing resource consumption. Furthermore, the soft prompting technique (Gao et al., 2023; Jia et al., 2022b) has demonstrated advancements in efficiently fine-tuning models, allowing them to achieve strong performance with language instruction (Wang et al., 2024b; Zhu et al., 2024), enabling effective execution of downstream tasks.
@@ 3.2.1. Rationales Generation with MLLM and LLM-judgment
Excellent rationales, serving as contextual semantic information, have been shown in many recent studies to enhance model learning (Lei et al., 2024c; Guo et al., 2025; Zhang et al., 2023; Zheng et al., 2023b). In this study, we distill richer semantic knowledge from reliable rationales generated by large models during training to improve Ov-GSR performance, enabling better recognition of rare and unseen situations. This process eliminates rationales at inference, enhancing model efficiency and deployment ability.
@@
3:function Multi-round LLM-judgment and refinement($P_{gli}^{Pseudo}$, $P_{gaz}^{Pseudo}$, $s$)
4:     $rating\leftarrow$ LLM-Judge($P_{gli/gaz}^{Pseudo},s$)
@@
6:         $P_{gli+/gaz+}^{refined}\leftarrow$ Refine-Rationale($P_{{gli/gaz}}^{{Pseudo}},s$)
7:         $rating\leftarrow$ LLM-Judge($P_{{gli+/gaz+}}^{{refined}},s$)
@@
In this process, we draw inspiration from the “single answer grading” judgment method (Zheng et al., 2023a), where the LLM-judge directly assigns a score to a rationale for describing the situation $s^{b}$ as shown in Algorithm 1. We use the same LLM to refine the rationales through multiple rounds of judgment and refinement if the assigned score is low (e.g., $<N=8$), ensuring that the generated rationales accurately describe the depicted situation. This process facilitates the accurate generation of positive glimpse $\mathbf{P}_{gli+}$ and gaze $\mathbf{P}_{gaz+}$ rationales. Additionally, we introduce a step where the LLM produces negative rationales $\mathbf{P}_{gli-}$ and $\mathbf{P}_{gaz-}$ by leveraging the general outputs of the MLLM. The negative rationale generation stays closely aligned with the positive text features but introduces variations in attribute information, helping the model better distinguish unseen and rare situations with negative distance loss (Eq. 3). These accurately refined rationales from large models (Guo et al., 2025; Team et al., 2024) provide rich semantic information, enabling the student model to acquire generalization knowledge (Zhao et al., 2024b; Cui et al., 2024; Lei et al., 2024c) during distillation process, thereby enhancing its situation recognition ability.
@@ 3.2.2. Multimodal Interactive Prompt Distillation Framework
Table 1. Results (%) of Ov-GSR methods on the Ov-SWiG dataset, including three settings and five metrics evaluated on the dev and test set. The higher the number the better the performance. * denotes the model uses open-vocabulary settings (Wang et al., 2022; Ning et al., 2023; Lei et al., 2024c). Bold number represents highest accuracy.
@@ 3.3. Overall Training and Inference Process
During training, we generate text embeddings $t^{v}$ and $t^{r}$ using a frozen CLIP (Radford et al., 2021) text encoder for activities and entities that correspond to the role classes in the situation set (Wang et al., 2022; Lei et al., 2024c). Then, we compute the similarity between activity text embeddings and visual features using matrix multiplication, where the activity visual embeddings are defined as $\mathbf{\varepsilon}_{s}^{v}=\beta(\mathbf{X}_{S}W^{v}\cdot t^{v})$. Here, $\mathbf{X}_{S}\in\mathbb{R}^{D\times H\times W}$ denotes the projected visual features from the frozen student model’s CLIP vision encoder (Radford et al., 2021), and $W^{v}$ is an additional projection layer serves as the activity head. The softmax function $\beta$ is applied for activity prediction. Similarly, we calculate the similarity between entity embeddings with the role visual embeddings $\mathbf{\varepsilon}_{s}^{r}=\beta(\mathbf{X}_{S}^{r}\mathbf{W}^{r}\cdot t^{r})$ with the additional multihead self-attention modules $\phi(\cdot)$ and projector as the role head, where $\mathbf{X}_{s}^{r}=\phi([\mathbf{X}_{S},\mathbf{\bar{X}}^{v}_{S}]W)$. $\mathbf{\bar{X}}^{v}_{S}$ represents the mean-pooled activity features, which are added to guide the prediction of roles. This is based on the constraint that a situation is only considered correct if the activity is accurately predicted for GSR (Pratt et al., 2020; Cheng et al., 2022). The classification loss for the situation objective is:
@@ 4.1. Experimental Settings
Benchmark dataset. We evaluate our Ov-GSR approach on the newly split Ov-SWiG dataset, built upon SWiG (Pratt et al., 2020), containing 124,384 images split into 73,984 for training, and 25,200 each for development and test sets. Each image is paired with three verb frames annotated by three different annotators. We use 500 images within dev and test set to evaluate open-vocabulary performance for 1,500 unseen situation pairs, where 10 unseen verbs are randomly picked from frequently used to rarely used for this evaluate. This resulting 67 entity annotation is not seen in the training set. We select the last 20 rarely seen verbs, resulting in 3,000 rare situation pairs to evaluate the rare cases. Ov-SWIG dataset has 504 verb categories, 190 semantic role types, and 9928 object categories, where each verb is associated with 1 to 6 semantic roles. In addition, we further conduct experiments on the relevant HICO-DET dataset (Chao et al., 2018). It consists of 600 interaction combinations, encompassing 117 human actions and 80 objects. Following (Lei et al., 2024c; Wang et al., 2022), we simulate a zero-shot detection setting by excluding 120 rare interactions from the full set of 600.
Evaluation metrics. We follow prior GSR works (Pratt et al., 2020; Cheng et al., 2022; Cho et al., 2022) and adopt five evaluation metrics to assess our method: (1) verb: activity prediction accuracy; (2) value: entity prediction accuracy per role; (3) val-all: entity prediction accuracy for the full semantic role set; (4) grnd: grounding (localization) accuracy per role; and (5) grnd-all: grounding accuracy across all semantic role set. A grounding prediction is considered correct if its Intersection-over-Union (IoU) with the ground truth is $\geq$ 0.5. Metrics are evaluated under three settings: Top-1-all, Top-1-rare, and Top-1-unseen. Semantic role predictions (e.g., val, grnd) are deemed incorrect if the activity prediction is incorrect. We evaluate the performance of HICO-DET using mean Average Precision (mAP) that is the same as the existing methods  (Chao et al., 2018; Lei et al., 2024c). An HOI triplet prediction is a true positive if the IoU between both the human and object bounding boxes exceeds 0.5, and the predicted interaction category is correct.
Implementation Details. We utilize frozen CLIP (Radford et al., 2021) (CLIP-ViT-L14) as the student model and InstrutBlip (Dai et al., 2023) as the MLLM teacher to conduct experiment. All the dimensions of visual and text embeddings will project to D=512 in the experiment. The training learning rate of the proposed model is $10^{-4}$. We use AdamW Optimizer] with a weight decay of $10^{-4}$, where $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We train our proposed model for 10 epochs with a batch size of 32 on a single RTX3090 GPU, including the model analysis on figure 2.
@@ 4.2. Comparisons with existing methods
Table 1 compares the performance of Multimodal Interactive Prompt Distillation (MIPD) on the Ov-SWIG benchmark with existing approaches on the dev and test set. These approaches include the grounded situation recognition methods OpenSU (Liu et al., 2023b) and ClipSiTU (Roy et al., 2024) with an open-vocabulary classifier(Ning et al., 2023; Lei et al., 2024c). Additionally, we compare MIPD with open-vocabulary HOI methods trained end-to-end under the same experimental settings, such as THID (Wang et al., 2022) and CMD-SE (Lei et al., 2024c), which leverage knowledge from large foundation models. These methods are re-implemented using a same setting to ours for a standardized and fair comparison. The 190 semantic roles (Pratt et al., 2020) are seen all the time, so the value-all and grand-all matrices for roles are not included in rare and unseen scenarios. We observe that existing GSR models (Liu et al., 2023b; Roy et al., 2024) primarily improve Top-1-all performance but struggle with rare and unseen cases, suggesting a prediction bias toward frequent and previously seen situations. In contrast, open-vocabulary methods (Wang et al., 2022; Lei et al., 2024c) demonstrate better performance in rare and unseen cases. However, our proposed MIPD outperforms existing methods across top-1-all, rare, and unseen prediction settings. It improves unseen recognition and better identifies rare situations compared to previous approaches, demonstrating its strong generalization capabilities. We further apply our method to HOI detection and present the performance in Table 2, where our proposed approach outperforms other open-vocabulary HOI methods (Wang et al., 2022; Lei et al., 2024c) under the same end-to-end training setting (without an additional pre-trained object detector (Carion et al., 2020)) with ViT-B16, achieving superior unseen performance of 17.84% and rarely seen performance of 25.45%, demonstrating its effectiveness. Moreover, we follow the setting of ClipSitu (Roy et al., 2024) and compare the performance of the MIPD with existing GSR approaches in Table 3 on the closed-set SWiG dataset (Pratt et al., 2020), the result further validates the effectiveness of our proposed method. The improvement is partially attributed to better recognition of rare situations, as shown in Tables 1 and 2, which contributes to overall GSR performance gains. This further underscores the importance of distilling rich generalized multimodal knowledge from MLLMs to improve situation recognition.

### v2

@@ 1. Introduction
The ability to conduct situation recognition in the scene is one of the essential roles of vision (Dai et al., 2023; Liu et al., 2023a; Zhang et al., 2025; Lu et al., 2025) and language (Touvron et al., 2023; Li et al., 2023b) research, with broad applications in assistive technology such as autonomous driving (Pan et al., 2024; Liao et al., 2024; Wang et al., 2024a) and visual impairments (Zhao et al., 2024a). Recently, Multimodal Large Language Models (MLLMs) (Dai et al., 2023; Zhu et al., 2023a) have demonstrated remarkable zero-shot scene understanding and can be applied across various domains (Gao et al., 2025; Cai et al., 2024a; Li et al., 2023a). However, many of them rely on smaller large language model (LLM) counterparts (e.g., InstructBLIP (Dai et al., 2023), TinyLLaVA (Liu et al., 2024)), which often underperform in tasks like Grounded Situation Recognition (GSR) (Pratt et al., 2020; Jiang and Riloff, 2023) that require deep comprehension (Wei et al., 2022b; Lu et al., 2024), as reflected by the Top-1 activity prediction accuracy (Top-1-all-verb) in Figure 2. Moreover, although these models are relatively small compared to much larger ones (e.g., with 34B (Young et al., 2024; Touvron et al., 2023) parameters), fine-tuning and deploying such massive models for GSR remain challenging due to their substantial computational and resource demands. This issue is particularly critical for apply GSR to many assistive technologies, which often depend on low-resource edge devices rather than heavy servers with modern GPUs (Cai et al., 2024b; Wang et al., 2023; Mittal, 2024). Addressing this problem is crucial for advancing the development of small and efficient GSR model capable of accurately interpreting complex scenes while preserving the generalization capabilities of MLLMs, potentially benefiting a wide range of assistive technologies.
@@ 2.2. Knowledge Distillation of Large Models
Recent advanced works distill large model capabilities into smaller ones (Zhang et al., 2024; Zhao et al., 2024c; Sun et al., 2023; Wei et al., 2024), demonstrating promising results. Minmax (Wang et al., 2025) formulated dataset distillation as a minmax optimization problem and proposed neural characteristic function discrepancy to effectively measure distributional differences, enabling compact and high-quality synthetic dataset generation. PRR (Zhao et al., 2024c) introduces a retrieval-based Chain-of-Thought (CoT) (Wei et al., 2022c) distillation technique, which transfer knowledge from LLMs to smaller language models, enhancing the performance of the question answering tasks. Tinyllm (Tian et al., 2024) introduces a new knowledge distillation paradigm, where a small student LLM learns from multiple large teacher models, effectively capturing knowledge from multiple rationals while maintaining efficiency. Some methods focus on knowledge distillation through efficient prompt-tuning (Jia et al., 2022; Liu et al., 2025; Bahng et al., 2022; Chen et al., 2023). PromptMM (Wei et al., 2024) enhances recommender systems by leveraging prompt-tuning, enabling efficient distillation to bridge the semantic gap across multi-modal contexts. PromptKD (Li et al., 2024) leverages soft prompt-based imitation on unlabeled domain images, allowing a lightweight target model to acquire knowledge from a large teacher model through a novel unsupervised distillation approach. Differently, this work encapsulates the information of rich rationales and visual prompts from the teacher model to effectively distill multimodal knowledge into a student model for improved Ov-GSR.
@@ 2.3. Open-vocabulary Tasks
The most similar work to Ov-GSR is end-to-end open-vocabulary HOI detection (Wang et al., 2022; Lei et al., 2024b), which simultaneously recognizes actions in an image while detecting humans and objects. THID (Wang et al., 2022) distills and utilizes transferable knowledge from the pretrained CLIP model, integrating multimodal features into a joint visual-text space to enhance open-vocabulary interaction detection. CMD-SE (Lei et al., 2024b) presents a novel HOI detection framework that distills fine-grained human body part semantic knowledge from LLM to enhance interaction recognition. The proposed Ov-GSR focuses on recognizing the activity first, followed by the detection of multiple entities within an image. In contrast to CMD-SE, our model adopts a different approach by aligning multimodal knowledge from semantics, scene to instance-level, effectively distilling this knowledge to bridge the gap between seen and unseen situations while enhancing rare situation awareness.
@@ 3.1. Problem Overview and Motivation
Distill Knowledge from Large Models with Prompts: Knowledge distillation (Xu et al., 2024; Gu et al., 2024a; Ning et al., 2023) has emerged as a key technique to alleviate the substantial computational demands of modern MLLMs by training smaller student models to replicate the behavior of larger teacher models, significantly reducing resource consumption. Furthermore, the soft prompting technique (Gao et al., 2023; Jia et al., 2022) has demonstrated advancements in efficiently fine-tuning models, allowing them to achieve strong performance with language instruction (Wang et al., 2024b; Zhu et al., 2024), enabling effective execution of downstream tasks.
@@ 3.2.1. Rationales Generation with MLLM and LLM-judgment
Excellent rationales, serving as contextual semantic information, have been shown in many recent studies to enhance model learning (Lei et al., 2024b; Guo et al., 2025; Zhang et al., 2023; Zheng et al., 2023b). In this study, we distill richer semantic knowledge from reliable rationales generated by large models during training to improve Ov-GSR performance, enabling better recognition of rare and unseen situations. This process eliminates rationales at inference, enhancing model efficiency and deployment ability.
@@
[…]
@@
[…]
@@
In this process, we draw inspiration from the “single answer grading” judgment method (Zheng et al., 2023a), where the LLM-judge directly assigns a score to a rationale for describing the situation $s^{b}$ as shown in Algorithm 1. We use the same LLM to refine the rationales through multiple rounds of judgment and refinement if the assigned score is low (e.g., $<N=8$), ensuring that the generated rationales accurately describe the depicted situation. This process facilitates the accurate generation of positive glimpse $\mathbf{P}_{gli+}$ and gaze $\mathbf{P}_{gaz+}$ rationales. Additionally, we introduce a step where the LLM produces negative rationales $\mathbf{P}_{gli-}$ and $\mathbf{P}_{gaz-}$ by leveraging the general outputs of the MLLM. The negative rationale generation stays closely aligned with the positive text features but introduces variations in attribute information, helping the model better distinguish unseen and rare situations with negative distance loss (Eq. 3). These accurately refined rationales from large models (Guo et al., 2025; Team et al., 2024) provide rich semantic information, enabling the student model to acquire generalization knowledge (Zhao et al., 2024b; Cui et al., 2024; Lei et al., 2024b) during distillation process, thereby enhancing its situation recognition ability.
@@ 3.2.2. Multimodal Interactive Prompt Distillation Framework
Table 1. Results (%) of Ov-GSR methods on the Ov-SWiG dataset, including three settings and five metrics evaluated on the dev and test set. The higher the number the better the performance. * denotes the model uses open-vocabulary settings (Wang et al., 2022; Ning et al., 2023; Lei et al., 2024b). Bold number represents highest accuracy.
@@ 3.3. Overall Training and Inference Process
During training, we generate text embeddings $t^{v}$ and $t^{r}$ using a frozen CLIP (Radford et al., 2021) text encoder for activities and entities that correspond to the role classes in the situation set (Wang et al., 2022; Lei et al., 2024b). Then, we compute the similarity between activity text embeddings and visual features using matrix multiplication, where the activity visual embeddings are defined as $\mathbf{\varepsilon}_{s}^{v}=\beta(\mathbf{X}_{S}W^{v}\cdot t^{v})$. Here, $\mathbf{X}_{S}\in\mathbb{R}^{D\times H\times W}$ denotes the projected visual features from the frozen student model’s CLIP vision encoder (Radford et al., 2021), and $W^{v}$ is an additional projection layer serves as the activity head. The softmax function $\beta$ is applied for activity prediction. Similarly, we calculate the similarity between entity embeddings with the role visual embeddings $\mathbf{\varepsilon}_{s}^{r}=\beta(\mathbf{X}_{S}^{r}\mathbf{W}^{r}\cdot t^{r})$ with the additional multihead self-attention modules $\phi(\cdot)$ and projector as the role head, where $\mathbf{X}_{s}^{r}=\phi([\mathbf{X}_{S},\mathbf{\bar{X}}^{v}_{S}]W)$. $\mathbf{\bar{X}}^{v}_{S}$ represents the mean-pooled activity features, which are added to guide the prediction of roles. This is based on the constraint that a situation is only considered correct if the activity is accurately predicted for GSR (Pratt et al., 2020; Cheng et al., 2022). The classification loss for the situation objective is:
@@ 4.1. Experimental Settings
Benchmark dataset. We evaluate our Ov-GSR approach on the newly split Ov-SWiG dataset, built upon SWiG (Pratt et al., 2020), containing 124,384 images split into 73,984 for training, and 25,200 each for development and test sets. Each image is paired with three verb frames annotated by three different annotators. We use 500 images within dev and test set to evaluate open-vocabulary performance for 1,500 unseen situation pairs, where 10 unseen verbs are randomly picked from frequently used to rarely used for this evaluate. This resulting 67 entity annotation is not seen in the training set. We select the last 20 rarely seen verbs, resulting in 3,000 rare situation pairs to evaluate the rare cases. Ov-SWIG dataset has 504 verb categories, 190 semantic role types, and 9928 object categories, where each verb is associated with 1 to 6 semantic roles. In addition, we further conduct experiments on the relevant HICO-DET dataset (Chao et al., 2018). It consists of 600 interaction combinations, encompassing 117 human actions and 80 objects. Following (Lei et al., 2024b; Wang et al., 2022), we simulate a zero-shot detection setting by excluding 120 rare interactions from the full set of 600.
Evaluation metrics. We follow prior GSR works (Pratt et al., 2020; Cheng et al., 2022; Cho et al., 2022) and adopt five evaluation metrics to assess our method: (1) verb: activity prediction accuracy; (2) value: entity prediction accuracy per role; (3) val-all: entity prediction accuracy for the full semantic role set; (4) grnd: grounding (localization) accuracy per role; and (5) grnd-all: grounding accuracy across all semantic role set. A grounding prediction is considered correct if its Intersection-over-Union (IoU) with the ground truth is $\geq$ 0.5. Metrics are evaluated under three settings: Top-1-all, Top-1-rare, and Top-1-unseen. Semantic role predictions (e.g., val, grnd) are deemed incorrect if the activity prediction is incorrect. We evaluate the performance of HICO-DET using mean Average Precision (mAP) that is the same as the existing methods  (Chao et al., 2018; Lei et al., 2024b). An HOI triplet prediction is a true positive if the IoU between both the human and object bounding boxes exceeds 0.5, and the predicted interaction category is correct.
Implementation Details. We utilize frozen CLIP (Radford et al., 2021) (CLIP-ViT-L14) as the student model and InstrutBlip (Dai et al., 2023) as the MLLM teacher to conduct experiment. All the dimensions of visual and text embeddings will project to D=512 in the experiment. The training learning rate of the proposed model is $10^{-4}$. We use AdamW Optimizer with a weight decay of $10^{-4}$, where $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We train our proposed model for 10 epochs with a batch size of 32 on a single RTX3090 GPU, including the model analysis on figure 2.
@@ 4.2. Comparisons with existing methods
Table 1 compares the performance of Multimodal Interactive Prompt Distillation (MIPD) on the Ov-SWIG benchmark with existing approaches on the dev and test set. These approaches include the grounded situation recognition methods OpenSU (Liu et al., 2023b) and ClipSiTU (Roy et al., 2024) with an open-vocabulary classifier (Ning et al., 2023; Lei et al., 2024b). Additionally, we compare MIPD with open-vocabulary HOI methods trained end-to-end under the same experimental settings, such as THID (Wang et al., 2022) and CMD-SE (Lei et al., 2024b), which leverage knowledge from large foundation models. These methods are re-implemented using a same setting to ours for a standardized and fair comparison. The 190 semantic roles (Pratt et al., 2020) are seen all the time, so the value-all and grand-all matrices for roles are not included in rare and unseen scenarios. We observe that existing GSR models (Liu et al., 2023b; Roy et al., 2024) primarily improve Top-1-all performance but struggle with rare and unseen cases, suggesting a prediction bias toward frequent and previously seen situations. In contrast, open-vocabulary methods (Wang et al., 2022; Lei et al., 2024b) demonstrate better performance in rare and unseen cases. However, our proposed MIPD outperforms existing methods across top-1-all, rare, and unseen prediction settings. It improves unseen recognition and better identifies rare situations compared to previous approaches, demonstrating its strong generalization capabilities. We further apply our method to HOI detection and present the performance in Table 2, where our proposed approach outperforms other open-vocabulary HOI methods (Wang et al., 2022; Lei et al., 2024b) under the same end-to-end training setting (without an additional pre-trained object detector (Carion et al., 2020)) with ViT-B16, achieving superior unseen performance of 17.84% and rarely seen performance of 25.45%, demonstrating its effectiveness. Moreover, we follow the setting of ClipSitu (Roy et al., 2024) and compare the performance of the MIPD with existing GSR approaches in Table 3 on the closed-set SWiG dataset (Pratt et al., 2020), the result further validates the effectiveness of our proposed method. The improvement is partially attributed to better recognition of rare situations, as shown in Tables 1 and 2, which contributes to overall GSR performance gains. This further underscores the importance of distilling rich generalized multimodal knowledge from MLLMs to improve situation recognition.